### v1

@@ 1 Introduction
However, collecting such trajectories is far from trivial. Existing task-driven methods, which rely on humans or machines executing pre-defined tasks, face the following limitations: human collection requires annotators to label entire trajectories and pre-define high-level tasks manually (Li et al., 2024; Lù et al., 2024), making it both costly and labor-intensive. Model-based synthesis also faces critical challenges: (1) it heavily depends on pre-defined high-level tasks (Lai et al., 2024), which not only limit the scalability of synthesized data but also constrain its diversity; and (2) it struggles to ensure data quality, as errors in intermediate steps or mismatched task objectives can lead to incomplete or incoherent trajectories (Murty et al., 2024b; Patel et al., 2024). Above mentioned issues pose a bottleneck for advancing GUI agents. These issues lead to a critical bottleneck for advancing GUI agents. Thus, effective trajectory construction methods are a clear desideratum for addressing these challenges.
@@ 2 Related Works
The recent proliferation of LLMs has significantly boosted researchers’ interest in developing language agents (Durante et al., 2024) to explore the digital world (Feng et al., 2024; Wu et al., 2024a). One line of work leverages the capabilities of fixed LLMs to create agents using methods like prompt engineering, model collaboration (Wu et al., 2023; Sun et al., 2023; Jia et al., 2024), code or tool use (Sun et al., 2024a), self-improvement (Shinn et al., 2024; Xu et al., 2024a; Cheng et al., 2024a), or integration with world or agent models (Hu and Shu, 2023; Jin et al., 2024; Zhang et al., 2023). Another line focuses on fine-tuning to augment models with agentic abilities, including (1) the ability to perceive the state of the computer, such as understanding screens (Cheng et al., 2024b; Gou et al., 2024; Wu et al., 2024b) or application UI trees (Xie et al., 2024; Zheng et al., 2024a), (2) the ability to generate actions (click, type, scroll, etc. Chen et al., 2024a), and (3) the flexibility to operate across diverse environments, including web (Yao et al., 2022; Deng et al., 2023), desktop (Kapoor et al., 2024; Niu et al., 2024), and mobile platforms (Li et al., 2024; Wang et al., 2024a). Collectively, these efforts pave the way for digital automation, with agents engaging across a diverse digital landscape.
@@
High-quality GUI data are essential for bridging models from the symbolic world (Xu et al., 2024b) to the digital world (Wu et al., 2024a), enabling the development of computer control agents. Such data empower models to propose plans, execute appropriate actions, and autonomously navigate diverse environments (Zeng et al., 2024; Pan et al., 2024). Rico (Deka et al., 2017) first introduces sequential GUI data for mobile apps, while MiniWob (Shi et al., 2017) provides low-level keyboard and mouse actions for web-based tasks. Since then, several works have expanded the availability of such data for mobile (Rawles et al., 2023; Zhang et al., 2024b; Lu et al., 2024; Chai et al., 2024), web (Liu et al., 2018; Lù et al., 2024; Murty et al., 2024a), and desktop (Chen et al., 2024a) applications. To effectively build computer agents, the best approach is to use trajectory data, which should consist of sequences containing GUI information, both low-level and high-level instructions, as well as corresponding actions (Li et al., 2024; Zhang et al., 2024a; Zheng et al., 2024b). However, acquiring such trajectories poses significant challenges. First, existing datasets often lack essential components. Second, the reliance on manual curation makes data collection costly and inefficient. Finally, current works are usually tailored to specific GUI (e.g., web-only), restricting their applicability in broader scenarios.
@@ 3.3 Trajectory Reward Model
The whole process is shown in Algorithm 1. By leveraging TRM, OS-Genesis ensures that synthesized trajectories are utilized effectively, allowing the training process to benefit from both high-quality data and diverse task scenarios.
@@ Appendix F Details of Trajectory Reward Model
The Trajectory Reward Model (TRM) primarily assesses the quality of agent trajectories by focusing on completion and coherence. Based on a high-level instruction to complete, the agent’s entire action history (e.g., low-level instructions), and screenshots from the last three timesteps, GPT-4o is prompted to assign a score between 1 and 5 for the trajectory. Instead of instruction and in-context learning (Sun et al., 2024b), we include in the prompt specific aspects of coherence and completion to consider, along with detailed descriptions of what each score from 1 to 5 represents. Given the similarity between mobile and web tasks, we apply the same TRM to both, as shown in prompt  21.

### v2

@@ 1 Introduction
However, collecting such trajectories is far from trivial. Existing task-driven methods, which rely on humans or machines executing pre-defined tasks, face the following limitations: human collection requires annotators to label entire trajectories and pre-define high-level tasks manually (Li et al., 2024; Lù et al., 2024), making it both costly and labor-intensive. Model-based synthesis also faces critical challenges: (1) it heavily depends on pre-defined high-level tasks (Lai et al., 2024), which not only limit the scalability of synthesized data but also constrain its diversity; and (2) it struggles to ensure data quality, as errors in intermediate steps or mismatched task objectives can lead to incomplete or incoherent trajectories (Murty et al., 2024b; Patel et al., 2024). Above mentioned issues pose a bottleneck for advancing GUI agents. Thus, effective trajectory construction methods are a clear desideratum for addressing these challenges.
@@ 2 Related Works
The recent proliferation of LLMs has significantly boosted researchers’ interest in developing language agents (Durante et al., 2024) to explore the digital world (Feng et al., 2024; Wu et al., 2024). One line of work leverages the capabilities of fixed LLMs to create agents using methods like prompt engineering, model collaboration (Wu et al., 2023; Sun et al., 2023b; Jia et al., 2024), code or tool use (Sun et al., 2024), self-improvement (Shinn et al., 2024; Xu et al., 2024a; Cheng et al., 2024a), or integration with world or agent models (Hu and Shu, 2023; Jin et al., 2024; Zhang et al., 2023). Another line focuses on fine-tuning to augment models with agentic abilities, including (1) the ability to perceive the state of the computer, such as understanding screens (Cheng et al., 2024b; Gou et al., 2024; Wu et al., 2025) or application UI trees (Xie et al., 2024; Zheng et al., 2024a), (2) the ability to generate actions (click, type, scroll, etc. Chen et al., 2024a), and (3) the flexibility to operate across diverse environments, including web (Yao et al., 2022; Deng et al., 2023), desktop (Kapoor et al., 2024; Niu et al., 2024), and mobile platforms (Li et al., 2024; Wang et al., 2024a). Collectively, these efforts pave the way for digital automation, with agents engaging across a diverse digital landscape.
@@
High-quality GUI data are essential for bridging models from the symbolic world (Xu et al., 2024b) to the digital world (Wu et al., 2024), enabling the development of computer control agents. Such data empower models to propose plans, execute appropriate actions, and autonomously navigate diverse environments (Zeng et al., 2024; Pan et al., 2024). Rico (Deka et al., 2017) first introduces sequential GUI data for mobile apps, while MiniWob (Shi et al., 2017) provides low-level keyboard and mouse actions for web-based tasks. Since then, several works have expanded the availability of such data for mobile (Rawles et al., 2023; Zhang et al., 2024b; Lu et al., 2024; Chai et al., 2024), web (Liu et al., 2018; Lù et al., 2024; Murty et al., 2024a), and desktop (Chen et al., 2024a) applications. To effectively build computer agents, the best approach is to use trajectory data, which should consist of sequences containing GUI information, both low-level and high-level instructions, as well as corresponding actions (Li et al., 2024; Zhang et al., 2024a; Zheng et al., 2024b). However, acquiring such trajectories poses significant challenges. First, existing datasets often lack essential components. Second, the reliance on manual curation makes data collection costly and inefficient. Finally, current works are usually tailored to specific GUI (e.g., web-only), restricting their applicability in broader scenarios.
@@ 3.3 Trajectory Reward Model
The whole process is shown in Algorithm 1. By leveraging TRM, OS-Genesis ensures that synthesized trajectories are utilized effectively, allowing the training process to benefit from both high-quality data and diverse task scenarios. Further details, along with consistency analyses between human annotators and different VLM backbones are shown in Appendix F.
@@ 6 Conclusion
Model-based Trajectory Construction.
We employ a model-based approach to build trajectories, thereby eliminating the need for human annotation. However, this also results in the quantity of successfully constructed trajectories being somewhat constrained by the capabilities of the task-executing model. We expect more advanced VLMs with GUI capability to address this in the future.
@@
Acknowledgement
This work is supported by Shanghai Artificial Intelligence Laboratory. We thank the reviewers of the SCI-FM Workshop @ ICLR 2025 and ACL Rolling Review for their valuable feedback, which has helped us improve this work.
Author Contributions
The authors contributed to this work in the following ways:
Project Leadership.
Qiushi Sun, Kanzhi Cheng, Zhiyong Wu.
OS-Genesis Concept.
Zhiyong Wu.
Data Curation.
Kanzhi Cheng, Qiushi Sun, Fangzhi Xu, Yian Wang, Zichen Ding, Liheng Chen, Chengyou Jia, Zhoumianze Liu, Chuanyang Jin.
Model and Training.
Zichen Ding, Kanzhi Cheng, Qiushi Sun, Zhenyu Wu.
Experiments and Analysis.
Qiushi Sun, Zichen Ding, Yian Wang, Kanzhi Cheng, Chuanyang Jin, Zhenyu Wu, Junxian He.
Paper Writing.
Qiushi Sun, Kanzhi Cheng, Zhiyong Wu, Junxian He, Chuanyang Jin, Zichen Ding.
Demos and Websites.
Qiushi Sun, Zhoumianze Liu.
Discussions.
All authors participate in research discussions and provide insightful technical advice.
Strategic Advice.
Junxian He, Guohao Li, Ben Kao, Yu Qiao.
@@ Appendix F Details of Trajectory Reward Model
The Trajectory Reward Model (TRM) primarily assesses the quality of agent trajectories by focusing on completion and coherence. Based on a high-level instruction to complete, the agent’s entire action history (e.g., low-level instructions), and screenshots from the last three timesteps, GPT-4o is prompted to assign a score between 1 and 5 for the trajectory. Instead of instruction and in-context learning (Sun et al., 2023a), we include in the prompt specific aspects of coherence and completion to consider, along with detailed descriptions of what each score from 1 to 5 represents. Given the similarity between mobile and web tasks, we apply the same TRM to both, as shown in prompt  21. Recognizing the critical role that TRM plays in ensuring data quality, additional analyses are conducted to validate its robustness.
F.1 Model Scoring Alignment
To demonstrate that models with varying capabilities can effectively serve as the TRM, we adopt Qwen2.5-VL-72B (Bai et al., 2025) as the TRM backbone and perform additional experiments on a small-scale dataset. Specifically, we sample 200 web and mobile trajectories for evaluation, results are shown in Table 5.
Domain Spearman Corr. p-val
Web 0.788 2.16e-22
Mobile 0.729 7.82e-18
Table 5: Consistency between GPT-4o and Qwen2.5-VL-72B as the backbone of TRM.
The results demonstrate that OS-Genesis can also work effectively with open-source models.
F.2 Human Scoring Alignment
To further validate the TRM, we sample 100 web and mobile trajectories and invite human annotators to label them according to the same guidelines used for the GPT-4o-based TRM. Subsequently, we calculated the Spearman correlation coefficients between the human annotations and the TRM-assigned rewards, as presented below.
Domain Spearman Corr. p-val
Mobile 0.813 9.23e-24
Web 0.798 2.54e-23
Table 6: Consistency between TRM and human annotators for reward modeling.
The results also demonstrate that TRM aligns well with human evaluations.